# A Neurosymbolic Fast and Slow Architecture for Graph Coloring

### v1

@@ Introduction
Graph coloring problems were generated using the Erdős–Rényi model, allowing for controlled variations in graph complexity through two main parameters:
Graph Size ($n$): Graphs varied $n\in[5,50]$ vertices, addressing a broad range of problem sizes.
@@
Edge Formation Probability ($p$): Set with $p\in[0.1,0.9]$, this parameter adjusted the density of edges, thereby influencing the difficulty of the coloring problem.
@@
Solvability Mix ($m$): The chromatic number for each graph, crucial for defining solvability, was calculated using the DSATUR algorithm. Based on this, problems were categorized into different solvability mixes, represented by the tuple $m=(a,b)$, where $a$ is the percentage of solvable problems and $b$ is the percentage of unsolvable problems:
  • $m=(100,0)$: All problems are solvable within the given constraints.
  • $m=(50,50)$: A balanced mix with equal solvable and unsolvable scenarios.
  • $m=(0,100)$: All problems are unsolvable, with the allowed number of colors less than the chromatic number.
Graphs were saved in DIMACS format to ensure standardization and enhance reproducibility.
SOFAI-v2 Implementation for Graph Coloring
The SOFAI-v2 architecture integrates advanced components designed to solve graph coloring problems dynamically. Figures 8 and 4 show example prompts provided to LLM (S1) for solving graph coloring. Below, we detail the implementation of these components using formal definitions and the functions they perform within the architecture:
  • Episodic Memory Implementation: The episodic memory, denoted as $M$, stores historical instances of solved graph coloring problems and their solutions. This aids the S1 solver by providing access to a set of similar past problems, enhancing the adaptability and effectiveness of the solver:
    • Memory Structure: Each entry in $M$ consists of graph attributes such as $nodes$, $edges$, $edge\_density$, $chromatic\_number$, and the corresponding $solution$.
    • Similarity Assessment: A similarity function $\sigma(x,M)$ evaluates the resemblance of the current graph attributes $x$ to those stored in $M$, prioritizing instances where attribute deviations are minimal. The similarity function is defined as:
      $\sigma(x,M)=\{(x_{j},f_{\text{v2}}(x_{j})):x_{j}\in M,\,d(x,x_{j})<\alpha\}$
      where, $d(x,x\_j)=|attributes(x)-attributes(x\_j)|$,   $\alpha$ threshold.
  • Improvement Trend Evaluation: If consecutive iterations indicate no improvement in solving capabilities, or if S1 fails to solve a problem within a set limit of five iterations, the S2 solver is invoked to ensure problem resolution.
  • Example Generation: The Graph Coloring Example Generator, $\mathcal{E}(x)$, simplifies the graph problem space by creating subgraphs. This module uses a greedy algorithm to propose potential solutions for these subgraphs, aiding the broader problem-solving strategy by focusing on simpler problem segments.
  • Feedback Generation and Validator: The Graph Coloring Validator assesses the validity of the proposed solutions and identifies errors such as coloring conflicts or uncolored vertices. Feedback denoted as $\mathcal{F}(f_{\text{S1}}(x,M_{x}))$ is provided based on these assessments, critical for the iterative adjustment of the solver’s strategy. The feedback process is formally incorporated as part of the metacognitive governance’s evaluation of $C(f_{\text{S1}}(x,M_{x}))$ against a confidence threshold $\theta_{\text{S1}}$.
This configuration ensures that SOFAI-v2 operates under a robust framework of real-time metacognitive governance, dynamically adapting to the complexity of graph coloring problems through experience-based learning and systematic feedback mechanisms. Examples of prompts, feedback types, and generated examples are in Supplementary Material.
@@
You are given an undirected graph with 3 colors available. Your task is to assign a color to each vertex such that no two adjacent vertices share the same color.
@@
Assign a unique color to each vertex, ensuring that no two vertices connected by an edge have the same color. Use no more than 3 distinct colors. Provide the color assignments for each vertex in the format:
@@
Please provide the color assignment for the new problem to solve directly below:
Figure 3: LLM prompt template for $m(100,0)$ for the graph coloring problem without using episodic memory.
SOFAI-v1 vs SOFAI-v2: SOFAI-v2 introduces significant advancements over its predecessor, SOFAI-v1, by incorporating iterative feedback mechanisms and episodic memory, which enhance the problem-solving capabilities of the S1 solver. Unlike SOFAI-v1, where problems not solved by S1 are directly escalated to S2 without iterative refinement, SOFAI-v2 systematically leverages historical data and adaptive feedback to optimize solutions before resorting to S2, thereby improving efficiency and efficacy.
@@
Figure 4: Example for episodic memory, which is added at the end of LLM prompt template in Fig 8.
Performance Metrics
All solver configurations, as detailed in Table 1, were evaluated using two primary metrics: success rate and average time taken. The success rate shown in Equation 1 measures the proportion of correct solutions out of the total number of problems, while the average time shown in Equation 2 taken calculates the mean time per problem solved within the time cap. The metrics are defined as follows:
@@
Average Time $\text{Average Time Taken}=\frac{\sum_{i=1}^{n}t_{i}}{n}$ (2)
where in Equation 2, $t_{i}$ represents the time taken to solve the $i$-th problem, and $n$ is the total number of problems solved within the time cap. Results were aggregated over three randomized trials to ensure reliability and consistency in the findings.
@@
SOFAI-v1 Combines S1 and S2 with rule-based metacognition
SOFAI-v2 Combines S1 and S2 with real-time metacognition
MC-S1-I[1-5] Iterative metacognitive feedback ranging from one to five iterations to refine S1 solutions.
Table 1: Descriptions of solver configurations employed in the experiments.
Results
This section presents an analysis of solver performance by systematically controlling key parameters. Unless mentioned, the edge formation probability is fixed to $p=0.5$, allowing a focused evaluation across varying graph sizes $n$ and solvability mixes $m$ to capture a broad range of problem configurations. Tables 2 and 3 provide the success rate and average time taken for each solver configuration under these controlled conditions. The Supplementary Material includes results for additional $p$ values to supplement this focused analysis, providing a comprehensive view of solver robustness across varying edge densities. From these results, we aim to address four key research questions (RQs) that provide deeper insights into SOFAI-v2 performance and robustness.
RQ 1 What improvement in success rate does SOFAI-v2 demonstrate compared to other solvers?
Graph Size m(100,0) (in %) m(0,100) (in %) m(50,50) (in %)
@@ Results
Table 2: Success rates (%) across solvers and problem configurations for graph sizes from 5 to 50 with edge probability $p=0.5$. Configurations $m=(100,0)$, $m=(0,100)$, and $m=(50,50)$ represent all solvable, all unsolvable, and a balanced mix of solvable and unsolvable instances, respectively.
According to the results shown in Table 2, SOFAI-v2, SOFAI-v1, and S2 exhibit comparable performance in fully solvable problem configuration ($m=(100,0)$). In contrast, in the unsolvable scenario ($m=(0,100)$) at a graph size of 50, SOFAI-v2 achieves a success rate of 53.85%, which is significantly higher than the 3.85% achieved by both SOFAI-v1 and S1. Similarly, in the mixed scenario ($m=(50,50)$), SOFAI-v2 continues to perform well with a 27.72% success rate, far surpassing SOFAI-v1 and S1, each at 2.11%.
The higher success rates of SOFAI-v2 in challenging problem configurations likely arise from its metacognitive governance that effectively identifies and resolves constraints, particularly in unsolvable configurations. This capability allows SOFAI-v2 to adapt more dynamically to complex problem sets, enhancing its performance in environments with a mix of solvable and unsolvable problems.
Answer to RQ 1: The SOFAI-v2 model shows enhanced performance in tackling unsolvable and mixed problem configurations, surpassing other solvers as detailed in Table 2. For example, within the unsolvable scenarios set (with $m=(0,100)$ at $n=50$), SOFAI-v2 achieves a success rate of 53.85%. This is approximately a 1298% improvement over the success rates of SOFAI-v1 and Solver S1, which are around 3.85%. The calculation for the improvement is derived as follows:
$\frac{53.85-3.85}{3.85}\times 100\%\approx 1298\%$
This significant performance boost is attributed to SOFAI-v2’s metacognitive governance system, which optimizes decision-making processes in complex problem configurations.
RQ 2 How does SOFAI-v2 compare in terms of average time taken for solving problems relative to other solvers?
Graph Size m(100,0) (in seconds) m(0,100) (in seconds) m(50,50) (in seconds)
5 3.23 0.001 3.23 0.92 3.13 0.001 3.13 0.77 3.23 0.001 3.23 0.92
10 4.56 0.19 4.69 2.21 4.08 0.17 4.14 1.07 5.28 0.22 5.48 3.74
15 5.61 70.85 55.46 53.23 6.20 72.02 39.89 22.33 5.12 67.08 70.01 85.37
20 6.79 199.27 154.11 120.86 7.52 198.70 111.00 46.96 6.68 196.22 193.15 195.25
30 11.53 200.00 137.80 117.10 10.87 200.00 82.30 52.10 9.92 200.00 192.85 193.67
40 9.69 200.00 153.41 144.54 9.46 200.00 109.71 101.18 10.30 200.00 194.47 193.67
50 41.98 200.00 200.00 139.17 62.76 200.00 88.93 88.93 21.83 200.00 200.00 194.47
@@
Table 3: Average time taken (in seconds) for different solvers across graph sizes and problem configurations, highlighting operational efficiency in solving or recognizing unsolvable problems.
According to the results in Table 3, SOFAI-v2 demonstrates distinct time efficiency advantages across various problem configurations and graph sizes. In fully solvable problems ($m=(100,0)$), the average solving time for SOFAI-v2 remains consistent and competitive with other solvers. However, in more complex problem configurations, such as all unsolvable problems ($m=(0,100)$) at $n=50$, SOFAI-v2 significantly reduces the average solving time to 88.93 seconds, a 55% decrease compared to S2, requiring up to 200 seconds. Similarly, in the mixed scenario ($m=(50,50)$) at the same graph size, SOFAI-v2 further improves efficiency, reducing the solving time to 194.47 seconds, demonstrating a nearly 3% improvement over SOFAI_v1 (200 seconds) and S2 (200 seconds).
The enhanced time efficiency of SOFAI-v2 in these challenging configurations likely results from its metacognitive governance, which optimizes computational strategies to handle complex and mixed problem configurations more efficiently. This capability enables SOFAI-v2 to solve problems faster and manage its computational resources more effectively, particularly in problem configurations where solvability varies widely.
Answer to RQ 2: In terms of time efficiency, SOFAI-v2 significantly outperforms other solvers, notably in problem configurations involving unsolvable and mixed problems. Specifically, at a graph size of 50 within the unsolvable configuration ($m=(0,100)$), the average time taken by SOFAI-v2 to attempt solving these problems is 88.93 seconds, which is approximately 55% less than the 200 seconds typically required by Solver S2. The percentage reduction is calculated as follows:
$\frac{200-88.93}{200}\times 100\%\approx 55\%$
This enhanced performance is attributed to the metacognitive governance within SOFAI-v2, which accelerates the problem-solving process and increases accuracy in complex problem configurations.
RQ 3 Does the iterative calling mechanism in SOFAI-v2 enhance the performance of its S1 solver through metacognitive governance?
@@
The iterative feedback mechanism within SOFAI-v2’s S1 solver shows a clear progression in performance across different graph sizes and problem configurations, as detailed in Figure 5. Each iteration, from MC-S1-I1 through MC-S1-I5, demonstrates a quantitative improvement in success rates, indicative of the solver’s enhanced problem-solving capabilities due to metacognitive governance.
In the fully solvable problems ($m=(100,0)$), the success rates of the initial iteration MC-S1-I1 start at approximately 20% and quickly improve to nearly 80% by MC-S1-I5 for $n=5$. This rapid improvement highlights the effectiveness of iterative feedback in optimizing the solver’s approach to solvable problems.
For the entirely unsolvable problems ($m=(0,100)$), success rates at $n=10$ begin at about 60% during MC-S1-I1 and rise to over 85% by MC-S1-I5. This increase of approximately 40% in success rate across iterations underscores the solver’s growing capability to efficiently recognize and handle unsolvable problem configuration without unnecessarily invoking higher-cost computational resources.
In the balanced mix problem configuration ($m=(50,50)$), iterative feedback leads to substantial improvement across graph sizes. Initially, success rates at MC-S1-I1 are around 40% for $n=10$ and increase to about 60% by MC-S1-I5. This demonstrates a consistent improvement of around 50% in discerning between solvable and unsolvable problems as iterations progress, illustrating the benefits of metacognitive governance in mixed problem configurations.
Answer to RQ 3: The iterative calling mechanism of SOFAI-v2 substantially enhances the performance of its S1 solver in various problem configurations. For instance, in the configuration where $m=(100,0)$, success rates escalate from 20% in the MC-S1-I1 to nearly 80% by the MC-S1-I5, marking an approximate 300% increase. This improvement is calculated as follows:
$\frac{80-20}{20}\times 100\%=300\%$
Similarly, in the configuration $m=(0,100)$, the success rates rise by 40%, from 60% in MC-S1-I1 to 85% in MC-S1-I5:
$\frac{85-60}{60}\times 100\%\approx 40\%$
In the balanced scenario ($m=(50,50)$), there is a 50% increase in success rates, from 40% in MC-S1-I1 to 60% in MC-S1-I5, demonstrating the solver’s adaptability:
$\frac{60-40}{40}\times 100\%=50\%$
The metacognitive governance within SOFAI-v2 facilitates these significant improvements across different configurations, enhancing the solver’s efficiency and adaptability.
RQ4 How does edge probability ($p$) influence the solvability and time efficiency of graph coloring problems across different solvers?
The impact of edge probability ($p$) on solvability and time efficiency across solvers was systematically evaluated through success rates and solving times as $p$ ranges from 0.1 to 0.9. As $p$ increases, all solvers generally exhibit a decreasing trend in success rates until reaching a plateau, as detailed in Figure 6. This trend indicates the increasing complexity of the graph coloring problems with more connections between nodes. Regarding time efficiency, there is a noticeable increase in average solving time, peaking before stabilizing, as shown in Figure 7.
SOFAI-v2 consistently outperforms S1, S2, and SOFAI-v1 in maintaining higher success rates across all edge probabilities, demonstrating enhanced resilience against increasing problem complexity. Additionally, SOFAI-v2 consistently requires less time to solve problems across all configurations, highlighting its superior time efficiency.
In the All Solvable ($m=(100,0)$) configuration, SOFAI-v2 exhibits comparable success rates as the probability $p$ increases, similar to other solvers. In the All Unsolvable ($m=(0,100)$) scenario, SOFAI-v2 achieves peak success rates at an edge probability of $p=0.1$, which are 20-40% higher than other solvers (SOFAI-v2 = 94%, SOFAI-v1 = 74%, Solver S1 = 68%, Solver S2 = 51%). This advantage is maintained with an approximate 10-20% lead as $p$ increases, as illustrated in Figure 6. Additionally, in this configuration, SOFAI-v2 solves problems 60-80% faster than its counterparts at mid-range probabilities ($p=0.1$ to $0.5$). For instance, at $p=0.1$, the solving times are as follows: SOFAI-v2 = 20 seconds, SOFAI-v1 = 55 seconds, and Solver S2 = 112 seconds, as depicted in Figure 7. In the Balanced Mix ($m=(50,50)$) scenario, the success rate for SOFAI-v2 peaks at approximately 72% at $p=0.1$, which is about 15-25% higher than the rates observed for SOFAI-v1, Solver S1, and Solver S2. The time efficiency in this scenario also notably surpasses other solvers, with reductions in solving times of 40-50%. Specifically, at $p=0.1$, the times are: SOFAI-v2 = 70 seconds, SOFAI-v1 = 100 seconds, and Solver S2 = 108 seconds.
Answer to RQ 4: Edge probability ($p$) substantially influences solvability and time efficiency, with SOFAI-v2 showing superior adaptability and operational efficiency. In problem configurations with increasing complexity, SOFAI-v2 sustains higher success rates and achieves significant time efficiency compared to other solvers, affirming its effective metacognitive governance.
Discussion
This study presented SOFAI-v2, a neurosymbolic architecture that integrates a fast, experience-based S1 powered by an LLM with a slow, deliberative S2 utilizing the DSATUR graph coloring algorithm governed by real-time metacognitive governance. Our experiments evaluated the performance of SOFAI-v2 on graph coloring problems with sizes ranging from 5 to 50, generated using the Erdős–Rényi model with an edge probability of $p=0.5$. We assessed the solvers across three problem configurations: all solvable ($m=(100,0)$), all unsolvable ($m=(0,100)$), and a balanced mix ($m=(50,50)$). The discussion addresses the research questions posed.
Integrating metacognitive governance and episodic memory in SOFAI-v2 demonstrates a notable enhancement in solving graph coloring problems, significantly improving success rates and reducing computational times across various graph sizes and problem complexities. Success rates in complex scenarios like the all-unsolvable configuration reveal SOFAI-v2’s superior adaptability, achieving up to a 53.85% success rate for larger graphs, a substantial improvement over SOFAI-v1 and traditional symbolic solvers. Additionally, iterative metacognitive feedback in SOFAI-v2’s System 1 has proven effective in refining problem-solving strategies, increasing efficiency, and enhancing the solver’s ability to differentiate between solvable and unsolvable problems. These capabilities underscore the value of combining neural and symbolic approaches to overcome limitations in traditional CSP solvers, demonstrating SOFAI-v2’s potential in dynamic and demanding CSP environments.
Implications and Comparisons with Prior Work

### v2

@@ 5.1 Data Generation and Problem Classification
Graph coloring instances were generated using the Erdős–Rényi model, controlling problem complexity through three primary parameters:
Graph Size ($n$): Varies between $n\in[5,50]$ vertices, encompassing a diverse spectrum of problem sizes.
@@
Edge Probability ($p$): Adjusted within $p\in[0.1,0.9]$, influencing edge density and problem difficulty.
@@
Solvability Mix ($m$): The chromatic number ($\chi(G)$) of each generated graph was computed using the DSATUR algorithm. Problems were then categorized based on solvability, represented as $m=(a,b)$, where $a$ and $b$ denote percentages of solvable and unsolvable instances respectively:
  • $m=(100,0)$: All problems solvable within constraints.
  • $m=(50,50)$: Equal distribution of solvable and unsolvable problems.
  • $m=(0,100)$: All problems unsolvable; set by choosing the number of available colors $k<\chi(G)$, ensuring infeasibility.
For each graph size and edge probability pair 100 graph problems were generated and stored in standardized DIMACS format for reproducibility. Github Repository: https://github.com/khvedant02/CSP-SOFAI_v2 .
5.2 SOFAI-v2 Implementation for Graph Coloring
SOFAI-v2 integrates iterative feedback mechanisms and episodic memory to dynamically resolve graph coloring problems. Figures 7 and 3 illustrate example prompts.
  • Episodic Memory ($M$): Stores historical graph instances with attributes—vertices, edges, edge density, chromatic number—and solutions. A similarity function $\sigma(x,M)$ retrieves relevant past solutions:
    $M_{x}=\{(x_{j},f_{\text{v2}}(x_{j})):x_{j}\in M,\,d(x,x_{j})<\alpha\}$
    where $d(x,x_{j})=|attributes(x)-attributes(x_{j})|$, and $\alpha$ is a similarity threshold.
  • Iterative Feedback Mechanism: If the solver’s correctness score (defined in Section 4) is insufficient, a two-step feedback process occurs:
    1. 1.
      Detect Conflict Pairs: Identify adjacent vertices assigned identical colors.
    2. 2.
      Template-Based Feedback: Algorithmically generate structured corrections (e.g., “Error: Vertices A and B are adjacent but have the same color”). Example of feedbacks is given in Supplementary Material Section 9.
  • Improvement Evaluation and S2 Invocation: If no solution improvement occurs over five iterations or if correctness remains below threshold, S2 is invoked as a fallback. If either solver identifies an instance as unsolvable, the instance is labeled as UNSAT. Exceeding the strict 200-second per-instance time limit results in failure, regardless of actual solvability.
  • Example Generation ($\mathcal{E}(x)$): A greedy-based algorithm generates simpler subgraph instances to guide S1. Examples are detailed further in Supplementary Material Section 9.
This comprehensive integration ensures SOFAI-v2 leverages both real-time metacognitive governance and experience-guided context for enhanced problem-solving effectiveness. Operationally, this mirrors a supervisory system that monitors conflict and recruits controlled processing when routine proposals fail (Norman & Shallice, 1986; Cooper & Shallice, 2000).
@@ 5.2 SOFAI-v2 Implementation for Graph Coloring
Please provide the color assignment for the new problem to solve, or respond with "NOT SOLVABLE" if it cannot be solved.
Figure 2: LLM prompt template for the graph coloring decision problem without using episodic memory.
@@
Figure 3: Example for episodic memory, which is added at the end of LLM prompt template in Supplementary Material Section 9.
SOFAI-v1 vs SOFAI-v2: SOFAI-v2 introduces substantial advancements over previous works of SOFAI (SOFAI-v1), by incorporating iterative feedback mechanisms and episodic memory, which enhance the problem-solving capabilities of the S1 solver. Unlike SOFAI-v1, where problems not solved by S1 are directly escalated to S2 without iterative refinement, SOFAI-v2 systematically leverages historical data and adaptive feedback to optimize solutions before resorting to S2, thereby improving efficiency and efficacy. Supplementary Material includes exact prompts, feedback templates, and our DSATUR+BT pseudocode.
5.3 Performance Metrics and Evaluation
All solver configurations (Table 1) were evaluated under a strict 200-second time limit per instance. Solver performance was measured using:
@@ 5.3 Performance Metrics and Evaluation
$\text{Average Time Taken}=\frac{\sum_{i=1}^{n}t_{i}}{n}$ (2)
where $t_{i}$ is the solving time for the $i$-th successfully solved instance and $n$ the total instances solved within the time limit. Instances exceeding this limit are marked as failures, explaining occasional 0% success rates even when solvable theoretically. Results reported are averaged over three randomized trials for reliability. To better understand solver efficiency beyond strict time constraints, we conducted supplementary experiments without time limits, solely for comparative runtime analysis. From a cognitive perspective, increased success with iterative MC reflects benefits of supervisory control over routine responding under higher constraint densities (Norman & Shallice, 1986; Cooper & Shallice, 2000).
Table 1: Descriptions of solver configurations employed in experiments.
@@
SOFAI-v1 Combines S1 and S2 with rule-based MC
SOFAI-v2 Combines S1 and S2 with iterative MC
MC-S1-I[1-5] Metacognitive feedback (1–5 iterations) to refine S1 solutions
5.4 Results
This section presents an analysis of solver performance by systematically controlling key parameters. Unless mentioned, the edge formation probability is fixed to $p=0.5$, allowing a focused evaluation across varying graph sizes $n$ and solvability mixes $m$ to capture a broad range of problem configurations. Tables 2 and 3 provide the success rate and average time taken for each solver configuration under these controlled conditions. The Supplementary Material 11 & 12 includes results for additional $p$ values to supplement this focused analysis, providing a comprehensive view of solver robustness across varying edge densities. From these results, we aim to address four key research questions (RQs) that provide deeper insights into SOFAI-v2 performance and robustness.
RQ 1 How does SOFAI-v2 success rate compare to that of S1, S2, SOFAI-v1?
Table 2: Success rates (%) across solvers and problem configurations for graph sizes from 5 to 50 with edge probability $p=0.5$. Configurations $m=(100,0)$, $m=(0,100)$, and $m=(50,50)$ represent solvable, unsolvable, and a balanced mix of solvable and unsolvable instances, respectively.
Graph Size m(100,0) [%] m(0,100) [%] m(50,50) [%]
@@ 5.4 Results
Table 3: Average time (seconds) for SOFAI_v1, SOFAI_v2, and S2 across graph sizes and configurations $m(100,0)$, $m(0,100)$, and $m(50,50)$.
Graph Size $m(100,0)$ (s) $m(0,100)$ (s) $m(50,50)$ (s)
SOFAI_v1 SOFAI_v2 S2 SOFAI_v1 SOFAI_v2 S2 SOFAI_v1 SOFAI_v2 S2
5 2 1 0 2 1 0 2 1 0
10 4 2 0 4 2 0 4 3 0
15 50 40 51 32 19 59 52 47 54
20 87 65 82 50 20 97 87 84 107
30 323 306 320 403 299 400 353 263 350
40 5043.2 4512.3 4995.8 5006.9 3511.7 4989.4 5012.1 3524.6 4992.7
50 20087.3 18034.9 19950.5 20044.2 14102.7 19920.6 20051.8 14178.3 19965.0
Answer to RQ 1: SOFAI-v2 demonstrates substantial improvements in success rates over other solvers, particularly in challenging problem configurations involving unsolvable or mixed solvability instances. Table 2 shows that for clearly solvable scenarios ($m=(100,0)$), SOFAI-v2 matches the optimal performance of SOFAI-v1 and S2 across smaller graph sizes. However, its key advantage becomes apparent in more complex scenarios:
  • For unsolvable scenarios ($m=(0,100)$) at graph size 50, SOFAI-v2 achieves a success rate of 53.85%, far surpassing the performance of SOFAI-v1 and S1, which each attain only 3.85%. This translates to an improvement of approximately 1298%.
  • In mixed scenarios ($m=(50,50)$), SOFAI-v2 maintains a success rate of 27.72%, substantially higher compared to just 2.11% by SOFAI-v1 and S1.
These performance gains arise primarily from SOFAI-v2’s enhanced metacognitive governance, which systematically leverages iterative solver feedback and episodic memory. This adaptive capability allows SOFAI-v2 to dynamically resolve constraints and effectively handle complex or initially challenging instances that other solvers cannot reliably solve.
For the above results we set a fixed time limit of 200 seconds per problem. If a solver does not finish within this time, it is counted as a failure. The average time taken results for this setting is reported separately in the Supplementary Material Section 10.
RQ 2 How does SOFAI-v2 compare in terms of average time relative to SOFAI-v1 and S2?
Answer to RQ 2: SOFAI-v2 demonstrates substantial improvements in time efficiency compared to SOFAI_v1 and S2, especially for complex configurations involving unsolvable or mixed scenarios. As detailed in Table 3, in the unsolvable scenario ($m=(0,100)$) at graph size $n=50$, SOFAI-v2 requires approximately 14,000 seconds, whereas both SOFAI_v1 and S2 take around 20,000 seconds. This corresponds to a substantial 30% reduction in average solving time. Similarly, for mixed solvability scenarios ($m=(50,50)$) at the same graph size ($n=50$), SOFAI-v2 achieves the same notable reduction of 30%, requiring only about 14,000 seconds compared to 20,000 seconds for SOFAI_v1 and S2. In fully solvable scenarios ($m=(100,0)$), the improvements remain consistent. For instance, at graph size 50, SOFAI-v2 completes solving tasks in roughly 18,000 seconds, a 10% time reduction compared to SOFAI_v1 and S2 (each requiring about 20,000 seconds).
These time-efficiency gains are attributed to SOFAI-v2’s advanced metacognitive governance, which intelligently manages solver interactions and computational resources based on real-time feedback and memory-driven decision-making. This dynamic resource allocation enhances solver efficiency, especially in challenging problem settings.
To ensure a fair and meaningful comparison of solver speed, we conducted this experiment without imposing any time limit. This is appropriate because SOFAI-v1, SOFAI-v2, and S2 are all complete solvers—that is, they can eventually solve or correctly determine the unsolvability of a problem given enough time. In contrast, S1 alone does not satisfy this property, so it is excluded from this timing comparison. This no-time-limit evaluation allows us to isolate and understand the time-efficiency advantages introduced by metacognitive governance in SOFAI-v2.
RQ 3 Does increasing the number of iterations in the feedback loop lead to an increased success rate in SOFAI-v2?
@@
Answer to RQ 3: Yes, the iterative calling mechanism substantially enhances the performance of SOFAI-v2’s S1 solver, driven by effective metacognitive governance. As illustrated in Figure 4, repeated solver invocations with structured feedback consistently improve success rates across all problem configurations:
  • In fully solvable problems ($m=(100,0)$), the initial success rate at iteration MC-S1-I1 (around 20% for graph size $n=5$) improves substantially to nearly 80% by iteration MC-S1-I5, a notable increase of 300%.
  • For completely unsolvable instances ($m=(0,100)$) at $n=10$, success rates begin at approximately 60% (MC-S1-I1) and increase to 85% by MC-S1-I5, an improvement of 40%.
  • In balanced solvability scenarios ($m=(50,50)$) at graph size $n=10$, success rates also see meaningful improvements, increasing from approximately 40% (MC-S1-I1) to 60% (MC-S1-I5), representing a 50% improvement.
These consistent improvements underscore how SOFAI-v2’s metacognitive governance improves solver responses through iterative feedback, enhancing overall adaptability and efficiency across diverse problem settings. For the above results, we set a fixed time limit of 200 seconds per problem. If a solver does not finish within this time, it is counted as a failure.
RQ4 How does the density of graph influence success rate and time across the considered solvers?
Answer to RQ 4: Edge probability ($p$) substantially affects both solvability and solving time across graph coloring solvers, with SOFAI-v2 demonstrating superior adaptability and efficiency. Figure 5 shows that increasing $p$ generally reduces success rates due to higher complexity from additional edges. However, SOFAI-v2 consistently maintains higher success rates compared to other solvers:
  • In fully solvable problems ($m=(100,0)$), all solvers maintain high success rates as $p$ increases, with SOFAI-v2 performing comparably to S2 and SOFAI-v1.
  • In completely unsolvable scenarios ($m=(0,100)$), SOFAI-v2 notably outperforms others, particularly at lower edge probabilities. For instance, at $p=0.1$, SOFAI-v2 achieves a 94% success rate, considerably surpassing SOFAI-v1 (74%), S1 (68%), and S2 (51%). This advantage remains pronounced (10-20%) even as $p$ increases.
  • In mixed scenarios ($m=(50,50)$), SOFAI-v2 again excels, achieving a peak success rate of about 72% at $p=0.1$, approximately 15-25% higher than other solvers.
Regarding time efficiency (Figure 6), as $p$ increases, solving times rise initially and then stabilize. However, SOFAI-v2 consistently requires less time across all problem configurations:
  • In the unsolvable scenario ($m=(0,100)$), SOFAI-v2 solves problems substantially faster at lower and mid-range probabilities. For instance, at $p=0.1$, SOFAI-v2 averages 20 seconds per problem, far less than SOFAI-v1 (55 seconds) and S2 (112 seconds).
  • For mixed solvability scenarios ($m=(50,50)$), SOFAI-v2 again demonstrates superior efficiency, 40-50% faster. Specifically, at $p=0.1$, SOFAI-v2 solves problems in about 70 seconds, compared to SOFAI-v1’s 100 seconds and S2’s 108 seconds.
Overall, SOFAI-v2’s robust metacognitive governance enables it to adapt efficiently and effectively to varying complexities induced by changing edge probabilities, achieving higher success rates and superior time efficiency. For the above results we set a fixed time limit of 200 seconds per problem. If a solver does not finish within this time, it is counted as a failure.
5.5 Other Considerations
We also explicitly address concerns regarding context window limitations inherent in modern LLMs. Although iterative prompting in our method increases token consumption, our analyses confirm that token usage remains comfortably within the Mistral-7B model’s 32k-token context limit, even under extreme conditions (graph size 100, edge probability 0.9, and five iterations):
  • Graph Representation (DIMACS format): approximately 1500 tokens
  • Initial Prompt & Instructions: approximately 500 tokens
  • Metacognitive Feedback per iteration (×5): approximately 2500 tokens (500 tokens each)
  • Episodic Memory Retrieval per iteration (×5): approximately 5000 tokens (1000 tokens each)
  • LLM Response per iteration (×5): approximately 5000 tokens (1000 tokens each)
  • Accumulated Prompt per iteration (including previous responses, ×5): approximately 15,000 tokens (3000 tokens each iteration)
In total, even the worst-case scenario results in approximately 24,500 tokens, well within Mistral-7B’s 32k-token context window. Consequently, context-window overflow is not an issue in our experiments. Operationally, the metacognitive governance budgets information (problem, feedback, episodic examples) so that fast/automatic proposals remain verifiable within resource bounds, aligning with supervised control over routine processing (Norman & Shallice, 1986; Cooper & Shallice, 2000).
Additionally, to account for the inherent variability in LLM output across runs, we repeat all experiments involving SOFAI-v1, SOFAI-v2, and S1 three times using the same set of problem instances. For success rate evaluation, a problem is considered a success if the solver succeeds in any of the three runs. For average time taken, we report the minimum time observed across the three trials. This choice reflects a conservative estimation of each system’s upper potential in ideal conditions and avoids penalizing performance due to stochastic failures. Such a best-of-$n$ (here n=3n{=}3) strategy is commonly used in LLM evaluations to mitigate randomness and highlight consistent solver capabilities (Kang et al., 2025; Chow et al., 2024). Cognitively, this also controls for noise in routine responding while assessing whether supervisory interventions reliably reduce errors.
We also note that SOFAI-based solvers (and S1) can declare UNSAT without a constructive certificate, whereas S2’s UNSAT arises from explicit search. This asymmetry can bias results on mixes containing unsatisfiable instances. However, the systematic variation in success rates across graph sizes and densities indicates the behavior is not arbitrary and reflects nontrivial problem sensitivity. A full proof-producing pipeline for UNSAT is orthogonal to our current scope; for automatic and quick evaluation, we accept such declarations, and we flag this as a limitation to be addressed in future work
5.6 Implications and Comparisons with Prior Work